# Thermodynamics and Gravitational Signatures of Rotating Black Holes in the Generalized Extended Uncertainty Principle

### v1

@@ Abstract
We investigate the phenomenological implications of quantum gravity on rotating black holes within the framework of the Generalized Extended Uncertainty Principle (GEUP), which incorporates both minimal length (ultraviolet) and large-scale (infrared) corrections. Lacking a full non-perturbative formulation of quantum gravity, we adopt a metric-based approach. We construct a stationary, axisymmetric ansatz via the Newman-Janis algorithm to model the kinematic features of a rotating black hole subject to Generalized Extended Uncertainty Principle (GEUP) corrections. The thermodynamic analysis reveals that in the infrared-dominated regime, the Hawking temperature scales as $T_{H}\sim M^{-3}$, leading to a rapid cooling phase that significantly prolongs the lifetime of supermassive black holes. We derive the modified Teukolsky Master Equation for gravitational perturbations and demonstrate that the background geometry preserves the isospectrality between axial and polar modes. In the eikonal limit, the quasinormal mode (QNM) spectrum exhibits orthogonal shifts: the minimal length parameter $\beta$ induces a spectral blueshift and enhanced damping, while the large-scale parameter $\alpha$ induces a spectral redshift and suppressed damping. Finally, we constrain the theory using observational data from LIGO/Virgo and the Event Horizon Telescope. We establish that the shadow of M87* is approximately $10^{6}$ times more sensitive to large-scale corrections than Sgr A*, placing stringent bounds on the EUP parameter, while gravitational wave spectroscopy provides complementary constraints on the GUP sector.
@@ I Introduction
The paper is organized as follows: Section II establishes the modified commutation relations and derives the effective GEUP mass and metric. Section III analyzes the thermodynamic quantities, presenting the heat capacity and stability profiles. Section IV investigates the gravitational wave signatures via the QNM spectrum. In section V investigates the shadow constraints. Finally, Section VI summarizes our findings and discusses future outlooks.
@@ II The Metric: Rotating GEUP Black Hole
$f(r)=1-\frac{2G\mathcal{M}(r)}{r}.$ (4)
In this framework, the quantum gravity corrections are naturally incorporated into the effective mass $\mathcal{M}(r)$. This approach preserves the algebraic structure of the Schwarzschild solution while modifying the radial dependence of the gravitational potential to account for both short-distance (GUP) and large-distance (EUP) deformations.
@@
III Thermodynamic Properties
@@ III.1 Thermodynamic Variables and the First Law
Equation (13) reveals a crucial physical implication of the GEUP framework. In the EUP-dominated regime ($M\gg M_{Pl}$), the effective mass scales as $\mathcal{M}\sim M^{3}/L_{*}^{2}$ [14, 17]. Consequently, the temperature scales as $T_{H}\sim M^{-3}$, falling much more rapidly than the standard Schwarzschild scaling ($T_{H}\sim M^{-1}$). This implies that large-scale quantum correlations effectively ”cool” supermassive black holes, significantly extending their evaporation lifetime compared to General Relativistic predictions.
The Bekenstein-Hawking entropy is proportional to the horizon area $A_{H}$. Unlike approaches that modify the area law itself (e.g., via logarithmic corrections [18]), the Mureika formalism encapsulates the deformation within the geometry:
@@
This consistency confirms that the Mureika formalism preserves the Legendre transformation structure of black hole thermodynamics [14], provided the variations are taken with respect to the renormalized mass parameter.
The Hawking temperature is derived from the surface gravity $\kappa$ or heuristically via the uncertainty principle. In the GEUP framework, the uncertainty relation $\Delta x\Delta p\gtrsim\hbar(1+\alpha(\Delta x)^{2})$ implies a modification to the temperature-mass relation. Identifying the position uncertainty with the horizon radius $\Delta x\sim r_{+}\approx 2M$, the modified temperature is:
@@
III.2 Thermal Stability Analysis
@@
The sign of $C_{J}$ determines stability: $C_{J}<0$ indicates instability (typical of Schwarzschild black holes), while $C_{J}>0$ indicates stability.
The stability profiles illustrated in Figure 1 are constructed parametrically by treating the bare mass $M$ as the independent variable. The GEUP corrections introduce a non-linear dependence in the effective mass $\mathcal{M}(M)$. Consequently, the horizon radius $r_{+}$ and the heat capacity $C_{J}$ in (Eq. 19) are coupled:
$r_{+}=G\mathcal{M}+\sqrt{G^{2}\mathcal{M}^{2}-a^{2}},$ (20)
where $a=J/M$ is the specific angular momentum. This parametric system captures how the renormalization of $\mathcal{M}$ modifies the phase transition relative to the observable horizon size.
In the Schwarzschild limit ($a\to 0$), the expression simplifies to $C_{J}=-2\pi r_{+}^{2}\propto-\mathcal{M}^{2}$. However, to assess the physical stability, we must consider the specific heat with respect to the bare mass parameter $M$, defined as $C_{GEUP}=\frac{d\mathcal{M}}{dT_{H}}$. While the heat capacity is physically defined as the response of the ADM mass to temperature changes ($C=\partial\mathcal{M}/\partial T_{H}$), we evaluate this quantity parametrically using the bare mass $M$. Using the chain rule $\frac{d\mathcal{M}}{dT_{H}}=\frac{d\mathcal{M}}{dM}(\frac{dT_{H}}{dM})^{-1}$, we observe a potential stability shift.
Since $T_{H}\sim 1/\mathcal{M}$, we have $\frac{dT_{H}}{d\mathcal{M}}<0$. Stability requires the total derivative $\frac{dT_{H}}{dM}>0$ (temperature increases as mass increases). Taking the derivative of the effective mass (Eq. 2):
$\frac{d\mathcal{M}}{dM}=1-\frac{\beta_{0}M_{Pl}^{2}}{2M^{2}}+\frac{3\alpha_{0}G^{2}M^{2}}{L_{*}^{2}}.$ (21)
@@
In the EUP regime, the slope of the mass function steepens due to the $M^{2}/L_{*}^{2}$ term. As illustrated in Figure 1, while the geometric heat capacity $C_{J}$ (Eq. 19) formally retains the algebraic structure of GR, the physical cooling rate is suppressed. The enhanced entropy scaling suggests that GEUP black holes are more thermodynamically robust against fluctuations than their GR counterparts, potentially delaying the onset of the runaway evaporation phase typically associated with the final stages of black hole life [20]. The roots of the denominator in Eq. (19) correspond to Davies points, marking phase transitions between stable and unstable configurations; the GEUP parameters $\alpha$ and $\beta$ shift these critical points, altering the domain of stability in the $(M,a)$ parameter space.
III.3 Evaporation Lifetime
@@
The EUP modification acts as a ’cryogenic’ mechanism, suppressing the evaporation rate for large black holes. We note that this behavior contrasts with black holes in Anti-de Sitter (AdS) space, where the large-scale potential typically induces positive specific heat ($T\sim M^{1/3}$). Here, the EUP-induced renormalization preserves the negative specific heat of the system but drastically enhances the cooling rate ($T\sim M^{-3}$), thereby prolonging the semi-classical lifetime.
@@ IV Gravitational Perturbations
Having established the thermodynamic stability of the Rotating GEUP black hole, we now turn to its dynamical stability and gravitational wave signature by studying the propagation of gravitational perturbations. Since the GEUP metric (Eq. 8) is isometric to the Kerr metric under the mapping $M\to\mathcal{M}$, the spacetime remains algebraically special (Petrov Type D). Crucially, this algebraic classification relies on our treatment of the renormalized mass $\mathcal{M}$ as a global constant fixed by the source’s bare mass, rather than a running coupling function of the radial coordinate $\mathcal{M}(r)$. This assumption preserves the Type D character of the background geometry, thereby allowing us to employ the standard Newman-Penrose formalism [21] to decouple the perturbation equations—a procedure first developed by Teukolsky [22].
@@ IV.4 Constraints from LIGO/Virgo Detection GW150914
We can use this observational bound to constrain the GEUP parameters in the regime of stellar-mass black holes, which complements the supermassive black hole constraints derived from shadow observations. The fractional deviation of the real frequency derived in Eq. (37) is:
@@
1. GUP (Small-Scale) Constraint: Unlike the supermassive M87* ($M\sim 10^{9}M_{\odot}$), the remnant of GW150914 ($M\sim 62M_{\odot}$) is closer to the Planck scale by seven orders of magnitude. While still far from the quantum gravity regime, it provides a tighter bound on the minimum length parameter $\beta$ than shadow observations. Assuming the deviation is dominated by the GUP term ($\alpha\to 0$):
@@
2. EUP (Large-Scale) Constraint: For the Extended Uncertainty Principle, the correction scales as $M^{2}$. Comparing the sensitivity of GW150914 to M87*:
@@
V Observational Signatures: The Black Hole Shadow
The most direct strong-field test of the Rotating GEUP metric is the geometric shape of its shadow, measurable by very long baseline interferometry such as the Event Horizon Telescope (EHT) [34]. The shadow boundary corresponds to the apparent position of the photon capture radius (light ring) projected onto the observer’s sky.
V.1 Null Geodesics and Photon Capture
@@
where $\lambda$ is an affine parameter. Due to the symmetries of the spacetime, the action $S$ is separable:
@@
For photons, the rest mass is $\mu=0$. The motion is characterized by two conserved quantities: the energy $E$ and the axial angular momentum $L_{z}$. We introduce the Carter constant $\mathcal{Q}$, which arises from the separability of the angular and radial equations. It is convenient to define the dimensionless impact parameters:
$\xi=\frac{L_{z}}{E},\quad\eta=\frac{\mathcal{Q}}{E^{2}}.$ (48)
The radial equation of motion takes the form:
$\Sigma\frac{dr}{d\lambda}=\pm\sqrt{\mathcal{R}(r)},$ (49)
where the radial potential $\mathcal{R}(r)$ is given by:
@@
The boundary of the black hole shadow corresponds to the critical photon orbits—those that are unstable and exist at a constant radius $r_{p}$ before falling into the hole or escaping to infinity. These spherical photon orbits are defined by the conditions:
@@
Applying the conditions in Eq. (51) to the potential, we obtain the critical impact parameters $(\xi_{c},\eta_{c})$ as parametric functions of the orbital radius $r_{p}$:
@@
These equations describe the locus of the shadow boundary. The crucial modification introduced by the GEUP is that the critical parameters scale with $\mathcal{M}$ rather than the bare mass $M$.
V.2 Celestial Coordinates and Shadow Radius
To an observer located at infinity with inclination angle $\theta_{0}$ (typically $\theta_{0}=90^{\circ}$ for an equatorial observer), the shadow appears as a dark region in the celestial sky. The celestial coordinates $(\alpha,\beta)$ are related to the impact parameters by:
$\alpha$ $=-\frac{\xi_{c}}{\sin\theta_{0}},$ (54)
$\beta$ $=\pm\sqrt{\eta_{c}+a^{2}\cos^{2}\theta_{0}-\xi_{c}^{2}\cot^{2}\theta_{0}}.$ (55)
For an equatorial observer ($\sin\theta_{0}=1,\cot\theta_{0}=0$), these simplify to $\alpha=-\xi_{c}$ and $\beta=\pm\sqrt{\eta_{c}}$.
While the rotation parameter $a$ introduces a D-shape asymmetry, the average radius of the shadow can be approximated by the Schwarzschild limit ($a\to 0$) for the purpose of order-of-magnitude constraints. In this limit, the photon sphere is located at $r_{p}=3\mathcal{M}$. Substituting this into the impact parameters yields $\xi_{c}=0$ and $\eta_{c}=27\mathcal{M}^{2}$. The shadow is perfectly circular with radius:
$R_{sh}=\sqrt{\alpha^{2}+\beta^{2}}=\sqrt{\eta_{c}}=3\sqrt{3}\mathcal{M}.$ (56)
Substituting the explicit form of the GEUP mass $\mathcal{M}\approx M(1+\frac{\beta M_{Pl}^{2}}{2M^{2}}-\frac{\alpha M^{2}}{2M_{Pl}^{2}})$, the modified shadow radius is:
$R_{sh}^{\text{GEUP}}\approx 3\sqrt{3}M\left(1+\frac{\beta M_{Pl}^{2}}{2M^{2}}-\frac{\alpha M^{2}}{2M_{Pl}^{2}}\right).$ (57)
This derivation confirms that the EUP parameter $\alpha$ reduces the apparent size of the black hole shadow, while the GUP parameter $\beta$ increases it.
V.3 Constraints from EHT Observations
We now quantify the viability of the GEUP model by comparing the theoretical radius $R_{sh}^{\text{GEUP}}$ with the shadow radii inferred from EHT observations of M87* [34] and Sagittarius A* (Sgr A*) [35].
It is important to address the morphology of the observed shadows. While the EHT images exhibit a clear brightness asymmetry (a ”crescent” shape) due to relativistic Doppler beaming of the accretion flow, the underlying geometry of the shadow boundary remains highly circular. The EHT analysis bounds the deviation from circularity to $\Delta C\lesssim 10\%$ for M87* [34]. This suggests that the spin-induced ”D-shape” deformation is either small (indicating moderate spin) or obscured by the resolution limit. Consequently, the angle-averaged shadow radius serves as the most robust observable for constraining the effective mass corrections.
We define the fractional deviation $\delta$ of this radius from the General Relativity prediction ($R_{GR}=3\sqrt{3}M$):
$\delta=\frac{R_{sh}^{\text{GEUP}}-R_{GR}}{R_{GR}}.$ (58)
Substituting Eq. (57) and the definition of $\mathcal{M}$, the deviation becomes:
$\delta\approx\frac{3\sqrt{3}\mathcal{M}-3\sqrt{3}M}{3\sqrt{3}M}=\frac{\mathcal{M}}{M}-1=\frac{\beta_{0}M_{Pl}^{2}}{2M^{2}}+\alpha_{0}\left(\frac{GM}{L_{*}}\right)^{2}.$ (59)
For supermassive black holes ($M\gg M_{Pl}$), the GUP term (proportional to $M^{-2}$) is negligible. The deviation is dominated by the EUP term (proportional to $M^{2}$):
$\delta\approx\alpha_{0}\left(\frac{GM}{L_{*}}\right)^{2}.$ (60)
1. Computation for M87*: The EHT observations of M87* ($M\approx 6.5\times 10^{9}M_{\odot}$) constrain the shadow radius to be consistent with GR within a $1\sigma$ confidence interval of approximately $17\%$ [36]. Setting $|\delta|\leq 0.17$ in Eq. (60):
$\alpha_{0}\left(\frac{G\cdot 6.5\times 10^{9}M_{\odot}}{L_{*}}\right)^{2}\lesssim 0.17.$ (61)
Solving for the fundamental EUP coupling $\alpha_{0}$:
$\alpha_{0}\lesssim 0.17\left(\frac{L_{*}}{R_{M87}}\right)^{2},$ (62)
where $R_{M87}=GM_{M87}\approx 10^{10}$ km is the gravitational radius.
2. Computation for Sgr A*: The observations of Sgr A* ($M\approx 4.0\times 10^{6}M_{\odot}$) are consistent with the Schwarzschild prediction within roughly $10\%$ [37]. Setting $|\delta|\leq 0.10$:
$\alpha_{0}\left(\frac{G\cdot 4.0\times 10^{6}M_{\odot}}{L_{*}}\right)^{2}\lesssim 0.10.$ (63)
The ratio of the constraints reveals the sensitivity of the different targets. Comparing the EUP terms:
$\frac{\delta_{M87}}{\delta_{SgrA}}=\left(\frac{M_{M87}}{M_{SgrA}}\right)^{2}\approx\left(\frac{6.5\times 10^{9}}{4.0\times 10^{6}}\right)^{2}\approx 2.6\times 10^{6}.$ (64)
This calculation demonstrates that M87* is approximately a million times more sensitive to EUP-type (large scale) corrections than Sgr A*. A value of $\alpha_{0}$ that produces a detectable $10\%$ deviation in Sgr A* would produce a physically impossible $260,000\%$ deviation in M87*. Therefore, the null result from M87* ($|\delta|\lesssim 0.17$) places the definitive upper bound on the Extended Uncertainty Principle parameter space. The data confirms that for astrophysical black holes, the effective mass renormalization must remain perturbative, $\mathcal{M}\approx M$.
V.4 Observational Distinguishability and Scale Invariance Breaking
It is important to address the phenomenological interpretation of the effective mass $\mathcal{M}$. Since the derived metric is formally identical to the Kerr solution with $M$ replaced by $\mathcal{M}$, a single geometric observation (e.g., a shadow radius measurement) yields a result consistent with a Kerr black hole of mass $M_{obs}=\mathcal{M}$. An observer without prior knowledge of the ”bare” parameter $M$ might interpret the object as a standard GR black hole.
However, the GEUP introduces a fundamental violation of the scale invariance inherent in General Relativity. In standard GR, dimensionless observables such as the product of the mass and the ringdown frequency, $\mathfrak{Im}=M\omega_{R}$, are universal constants for a given spin parameter $a_{*}$. In the GEUP framework, this universality is broken. The renormalization factor $\Xi=\mathcal{M}/M$ is mass-dependent:
$\frac{\mathcal{M}}{M}\approx 1+\beta\frac{M_{Pl}^{2}}{M^{2}}-\alpha M^{2}.$ (65)
Consequently, the dimensionless product observed across a population of black holes will exhibit a mass-dependent drift. While a single source limits us to measuring $M_{obs}$, a population analysis comparing stellar-mass black holes (LIGO regime, GUP-dominated) against supermassive black holes (EHT regime, EUP-dominated) would reveal that they do not fit a single Kerr mass-scaling curve. The ”deviation” reported in this work is therefore a deviation from the universal scaling relations of General Relativity.

### v2

# Phenomenology of Rotating GEUP Black Holes

@@ Abstract
We investigate the phenomenological implications of quantum gravity on rotating black holes within the framework of the Generalized Extended Uncertainty Principle (GEUP), which incorporates both minimal length (ultraviolet) and large-scale (infrared) corrections. Lacking a full non-perturbative formulation of quantum gravity, we adopt a metric-based approach. We construct a stationary, axisymmetric ansatz via the Newman-Janis algorithm to model the kinematic features of a rotating black hole subject to Generalized Extended Uncertainty Principle (GEUP) corrections. The thermodynamic analysis reveals that in the infrared-dominated regime, the Hawking temperature scales as $T_{H}\sim M^{-3}$, leading to a rapid cooling phase that significantly prolongs the lifetime of supermassive black holes. We derive the modified Teukolsky Master Equation for gravitational perturbations and demonstrate that the background geometry preserves the isospectrality between axial and polar modes. In the eikonal limit, the quasinormal mode (QNM) spectrum exhibits orthogonal shifts: the minimal length parameter $\beta$ induces a spectral blueshift and enhanced damping, while the large-scale parameter $\alpha$ induces a spectral redshift and suppressed damping. We discuss how the thermodynamic evolution of the black hole breaks the geometric degeneracy that exists for static observers.
@@ I Introduction
The paper is organized as follows: Section II establishes the modified commutation relations and derives the effective GEUP mass and metric. Section III analyzes the thermodynamic quantities, presenting the heat capacity and stability profiles. Section IV investigates the gravitational wave signatures via the QNM spectrum. Section V investigates the observational consistency of the model and discusses how the thermodynamic evolution breaks the geometric degeneracy. Finally, Section VI summarizes our findings and discusses future outlooks.
@@ II The Metric: Rotating GEUP Black Hole
$f(r)=1-\frac{2G\mathcal{M}}{r}.$ (4)
In this framework, the quantum gravity corrections are naturally incorporated into the constant effective mass $\mathcal{M}$. This approach preserves the algebraic structure of the Schwarzschild solution while modifying the strength of the gravitational potential to account for both short-distance (GUP) and large-distance (EUP) deformations.
@@
It is important to note that because the effective mass $\mathcal{M}$ is treated as a global constant determined by the source parameters $(M,\alpha,\beta)$ rather than a coordinate-dependent function, the metric in Eq. (8) retains the vacuum character of the standard Kerr solution. Direct calculation of the Einstein tensor yields $G_{\mu\nu}=0$, implying that the leading-order components of the effective energy-momentum tensor vanish ($\rho=p_{i}=0$). Consequently, the Weak Energy Condition ($\rho\geq 0,\rho+p_{i}\geq 0$) is trivially satisfied throughout the spacetime, ensuring that the phenomenological modifications do not introduce unphysical matter distributions in the domain of outer communication.
@@
III Thermodynamic Properties and Cryogenic Evolution
@@ III.1 Thermodynamic Variables and the First Law
Equation (13) reveals a crucial physical implication of the GEUP framework. In the EUP-dominated regime ($M\gg M_{Pl}$), the effective mass scales as $\mathcal{M}\sim M^{3}/L_{*}^{2}$ [14]. Consequently, the temperature scales as $T_{H}\sim\mathcal{M}^{-1}\sim M^{-3}$, falling much more rapidly than the standard Schwarzschild scaling ($T_{H}\sim M^{-1}$). This implies that large-scale quantum correlations effectively ”cool” supermassive black holes, significantly extending their evaporation lifetime compared to General Relativistic predictions.
The Bekenstein-Hawking entropy is proportional to the horizon area $A_{H}$. Unlike approaches that modify the area law itself, the Mureika formalism encapsulates the deformation within the geometry:
@@
This consistency confirms that the formalism preserves the Legendre transformation structure of black hole thermodynamics provided the variations are taken with respect to the renormalized mass parameter.
III.2 Heuristic Derivation of Temperature Limits
The Hawking temperature can also be understood heuristically via the uncertainty principle. In the GEUP framework, the uncertainty relation $\Delta x\Delta p\gtrsim\hbar(1+\alpha(\Delta x)^{2})$ implies a modification to the temperature-mass relation. Identifying the position uncertainty with the horizon radius $\Delta x\sim r_{+}\approx 2M$, the modified temperature is:
@@ III.2 Heuristic Derivation of Temperature Limits
To visualize the distinct thermodynamic behavior of the GEUP black hole, we present the temperature-mass phase diagram in Figure 1. The plot compares the GEUP trajectory against the standard Schwarzschild behavior [17] and the Schwarzschild-AdS case [18]. In the intermediate mass regime, the curve follows the standard General Relativistic scaling ($T_{H}\propto M^{-1}$). However, significant deviations appear in the asymptotic limits. At the Planck scale (left), the GUP correction forces the temperature to reach a maximum and vanish as $M\to M_{min}$, preventing the final evaporation singularity. Conversely, in the infrared regime (right), the trajectory diverges sharply from both the standard case and the AdS case. While large AdS black holes become thermodynamically stable with positive specific heat ($T_{AdS}\propto M^{1/3}$), the GEUP black hole enters a ’super-cooling’ phase where $T_{GEUP}\propto M^{-3}$. This steep negative slope visually confirms the rapid suppression of Hawking radiation for supermassive sources.
Figure 1: Log-log phase diagram comparing the Hawking temperature evolution for three black hole models. The dashed black line represents standard General Relativity ($T\propto M^{-1}$). The dot-dashed blue line represents the Schwarzschild-AdS behavior, where large masses become hot and thermodynamically stable ($T\propto M^{1/3}$). The solid red line depicts the Rotating GEUP black hole. Key features include the maximum temperature peak at the Planck scale (GUP regime) and the rapid cooling phase at large scales (EUP regime) where $T\propto M^{-3}$, contrasting sharply with the AdS behavior.
III.3 Thermal Stability Analysis
@@
The sign of $C_{J}$ determines stability: $C_{J}<0$ indicates instability (typical of Schwarzschild black holes), while $C_{J}>0$ indicates stability. The roots of the denominator in Eq. (19) correspond to Davies points, marking phase transitions between stable and unstable configurations; the GEUP parameters $\alpha$ and $\beta$ shift these critical points, altering the domain of stability in the $(M,a)$ parameter space.
@@
III.4 Cryogenic Evolution and Lifetime Enhancement
@@
The EUP modification acts as a ‘cryogenic’ mechanism, suppressing the evaporation rate for large black holes. We note that this behavior contrasts with black holes in Anti-de Sitter (AdS) space, where the large-scale potential typically induces positive specific heat ($T\sim M^{1/3}$). Here, the EUP-induced renormalization preserves the negative specific heat of the system but drastically enhances the cooling rate ($T\sim M^{-3}$), thereby prolonging the semi-classical lifetime.
Physically, this rapid cooling behavior ($T_{H}\sim M^{-3}$) can be understood as a consequence of the macroscopic ‘fuzziness’ introduced by the Extended Uncertainty Principle in the infrared regime. Unlike standard General Relativity, where the horizon radius is strictly proportional to the bare mass ($r_{+}\propto M$), the GEUP correction implies that position uncertainty grows with the scale of the system. For supermassive black holes, this macroscopic uncertainty effectively delocalizes the horizon, causing the effective gravitational radius to inflate disproportionately ($\mathcal{M}\sim M^{3}$). This geometric inflation drastically reduces the surface gravity ($\kappa\propto 1/r_{+}$), driving the black hole into a ‘cryogenic’ state much faster than the standard Hawking scaling would predict.
@@ IV Gravitational Perturbations
Having established the thermodynamic stability of the Rotating GEUP black hole, we now turn to its dynamical stability and gravitational wave signature by studying the propagation of gravitational perturbations. Since the GEUP metric (Eq. 8) is isometric to the Kerr metric under the mapping $M\to\mathcal{M}$. Spacetime remains algebraically special (Petrov Type D). Crucially, this algebraic classification relies on our treatment of the renormalized mass $\mathcal{M}$ as a global constant fixed by the source’s bare mass, rather than a running coupling function of the radial coordinate $\mathcal{M}(r)$. This assumption preserves the Type D character of the background geometry, thereby allowing us to employ the standard Newman-Penrose formalism [20] to decouple the perturbation equations—a procedure first developed by Teukolsky [21].
@@ IV.3.1 QNM Frequency Shift via the Eikonal Limit
It is important to note that the frequencies derived in Eqs. 35 and 36 rely on the Eikonal (geometric optics) approximation, which is formally valid in the limit of large multipole numbers ($l\gg 1$). While this limit robustly captures the qualitative behavior of the spectral shifts—specifically the orthogonal directions of the UV and IR corrections—observational ringdown signals (such as those detected by LIGO/Virgo) are dominated by the fundamental quadrupole mode ($l=2$). For precision waveform modeling required for data analysis, higher-order WKB methods or numerical integration would be necessary to quantify corrections of order $1/l$.
@@ IV.4 Constraints from LIGO/Virgo Detection GW150914
We can use this observational bound to constrain the GEUP parameters in the regime of stellar-mass black holes, which complements the constraints derived from shadow observations for supermassive black holes. The fractional deviation of the real frequency derived in Eq. (35) is:
@@
Unlike the supermassive M87* ($M\sim 10^{9}M_{\odot}$), the remnant of GW150914 ($M\sim 62M_{\odot}$) is closer to the Planck scale by seven orders of magnitude. While still far from the quantum gravity regime, it provides a tighter bound on the minimum length parameter $\beta$ than shadow observations. Assuming the deviation is dominated by the GUP term ($\alpha\to 0$):
@@
For the Extended Uncertainty Principle, the correction scales as $M^{2}$. Comparing the sensitivity of GW150914 to M87*:
@@
V Observational Implications
In the previous sections, we established that the GEUP corrections can be encoded into a renormalized effective mass $\mathcal{M}$, yielding a metric that is formally identical to the Kerr solution. This raises a fundamental issue of observability, the Indistinguishability Problem.
Gravitational measurements, whether via the Event Horizon Telescope (shadow radius) or LIGO/Virgo (inspiral chirp mass), measure the ADM mass of the system, which in our framework corresponds to $\mathcal{M}$, not the bare mass $M$. Consequently, for a static observer, a GEUP black hole is geometrically indistinguishable from a standard General Relativistic black hole of mass $\mathcal{M}$. The structural parameters $\alpha$ (EUP) and $\beta$ (GUP) are absorbed into the mass definition. Therefore, static geometric observations cannot constrain these parameters without an independent, non-gravitational determination of the source’s bare corpuscular content.
However, while the geometry is degenerate, the dynamics are distinct. The degeneracy is broken by the thermodynamic evolution of the system. In General Relativity, the evaporation rate is determined by $\dot{\mathcal{M}}\propto-\mathcal{M}^{-2}$. In the GEUP framework, the modified temperature scaling derived in Eq. (20) implies a radically different evolution law. Below, we demonstrate that while current shadow observations are consistent with the GEUP metric (by definition of the observed mass), the true test of the theory lies in the temporal variation of the mass (evaporation rate).
V.1 Null Geodesics and Critical Orbits
@@
where $\lambda$ is an affine parameter. The action $S$ is separable:
@@
For photons ($\mu=0$), the motion is characterized by the conserved energy $E$ and axial angular momentum $L_{z}$. Introducing the Carter constant $\mathcal{Q}$ and the dimensionless impact parameters $\xi=L_{z}/E$ and $\eta=\mathcal{Q}/E^{2}$, the radial equation of motion becomes $\Sigma\frac{dr}{d\lambda}=\pm\sqrt{\mathcal{R}(r)}$, where the radial potential is:
@@
The boundary of the black hole shadow corresponds to the critical photon orbits—those that are unstable and exist at a constant radius $r_{p}$. These spherical photon orbits are defined by the conditions:
@@
Solving these simultaneously yields the critical impact parameters $(\xi_{c},\eta_{c})$ as parametric functions of the orbital radius $r_{p}$:
@@
These equations describe the locus of the shadow boundary. To an observer at infinity with inclination $\theta_{0}$, the shadow coordinates $(\alpha_{sh},\beta_{sh})$ are:
$\alpha_{sh}$ $=-\frac{\xi_{c}}{\sin\theta_{0}},$ (50)
$\beta_{sh}$ $=\pm\sqrt{\eta_{c}+a^{2}\cos^{2}\theta_{0}-\xi_{c}^{2}\cot^{2}\theta_{0}}.$ (51)
@@
The crucial modification introduced by the GEUP is that the critical parameters scale with $\mathcal{M}$ rather than the bare mass $M$.
V.2 Consistency with EHT Observations
In the Schwarzschild limit ($a\to 0$), which suffices for order-of-magnitude estimates, the photon sphere is located at $r_{p}=3\mathcal{M}$, yielding a perfectly circular shadow of radius:
$R_{sh}=3\sqrt{3}\mathcal{M}.$ (52)
The EHT observations of M87* measured a shadow radius consistent with a mass of $\mathcal{M}_{M87}\approx 6.5\times 10^{9}M_{\odot}$ [33]. In our framework, this measurement fixes the value of the effective mass $\mathcal{M}$. Because the GEUP metric preserves the Kerr null-geodesic structure, the theoretical prediction for the shadow radius is $R_{sh}\approx 5.2\mathcal{M}$ (incorporating low spin corrections). The agreement between theory and observation is therefore guaranteed by identifying the observed mass with $\mathcal{M}$.
The significance of this result is not to constrain $\alpha$, but to ensure phenomenological consistency. The GEUP modification does not introduce exotic shadow distortions (such as fractal edges or cusps) that would contradict EHT images. The black hole looks ‘normal,’ which is a requirement for any viable effective field theory of gravity given current data.
V.3 Breaking the Degeneracy: Dynamical Signatures
How, then, can the GEUP model be falsified? The distinguishability lies in the Mass-Temperature relation. An observer monitoring a black hole over a cosmological timeframe would observe a deviation in the evaporation rate.
For a standard Kerr black hole, the luminosity $L\propto\mathcal{M}^{-2}$. For the GEUP black hole in the infrared regime (large mass), we found in Section III that $T_{H}\propto M^{-3}$. Assuming the perturbative limit $\mathcal{M}\approx M$, this implies a luminosity scaling $L\propto\mathcal{M}^{-6}$.
This suppression of Hawking radiation means that GEUP black holes are ‘dimmer’ and longer-lived than their GR counterparts. While the static shadow size ($R_{sh}\propto\mathcal{M}$) is indistinguishable at a single epoch, the time-derivative of the shadow size ($\dot{R}_{sh}\propto\dot{\mathcal{M}}$) follows a different trajectory. Thus, the GEUP signature is a dynamical suppression of evaporation, preserving primordial black holes far longer than the standard Hawking prediction.